1) Nicole Franco buys a Jeep Wrangler for $31,445, paying 20% down or $6,289 in cash and obtaining an auto loan for the balance. If the loan is to be repaid in equal annual installments at the end of each of 5 years at an interest rate of 4.75%,

a. What is the amount of each annual payment?

b. Create an amortization table?

Can you show it in Microsoft Excel Please
Answer #1

let me know if you need any clarification..

ans a
Cost of jeep 31445
Less Down payment 6289
loan amount 25156
using excel formula we can compute the annual installment
FV 0
PV -25156
N 5
I 4.75%
therefore PMT= $5,770.31 =PMT(4.75%,5,-25156,0,0)
ans b Amortization table
i ii=i*4.75% iii iv=i+ii-iii
Year Beginning loan Interest EMI paid Closing loan balance
0 25156
1         25,156.00 1,194.91    5,770.31    20,580.60
2         20,580.60      977.58    5,770.31    15,787.88
3         15,787.88      749.92    5,770.31    10,767.50
4         10,767.50      511.46    5,770.31      5,508.65
5           5,508.65      261.66    5,770.31                  -